OpenStudy (anonymous):

5 wholes 5/8 divided by 1 whole 1/4 show work please

OpenStudy (anonymous):

turn them into improper fractions: (8*5+5)/8 and (1*4+1)/4 which equals: 45/8 and 5/4 to divide, multiply the reciprical: (45/8) * (4/5) cross reduce: (9/2) * (1/1) = 9/2 = 4 wholes 1/2